# Unlocking the Potential of Model Calibration in Federated Learning

### v2

@@ B.2 Additional Results of Calibration on FL
Calibration Method FedAvg McMahan et al. [2017] FedProx Sahu et al. [2018] Scaffold Karimireddy et al. [2019] FedDyn Acar et al. [2021] FedNova Wang et al. [2020b]
Uncal. 92.18 1.96 1.04 91.76 3.21 0.96 91.50 5.03 1.32 97.74 3.28 0.50 93.80 1.95 0.69
Focal Lin et al. [2017] 90.75 9.02 2.28 90.76 10.93 2.35 88.74 16.49 5.38 97.53 2.38 0.47 93.50 8.32 1.82
LS Müller et al. [2019] 95.01 42.28 8.53 92.76 34.48 6.92 92.34 40.26 8.14 98.46 25.86 5.20 94.55 28.00 5.76
BS Brier [1950] 87.59 5.06 1.89 89.00 8.58 1.88 89.56 14.45 3.03 96.76 3.20 0.53 89.09 3.23 1.69
MMCE Kumar et al. [2018] 95.72 9.20 1.86 91.48 13.28 2.67 89.37 5.49 3.49 96.86 4.95 0.54 97.38 7.89 1.63
FLSD Mukhoti et al. [2020] 87.07 19.26 4.59 89.15 24.00 4.90 89.87 26.66 5.45 97.87 1.67 0.37 91.63 19.97 4.01
MbLS Liu et al. [2021] 93.48 14.22 3.08 91.97 12.60 2.76 89.32 8.29 4.58 98.02 6.16 1.32 94.34 9.21 2.12
DCA Liang et al. [2020] 92.89 1.06 0.81 92.45 1.62 0.70 92.34 3.49 1.12 97.75 2.18 0.41 94.34 0.87 0.49
MDCA Hebbalaguppe et al. [2022b] 92.44 1.93 0.98 91.81 2.78 0.91 92.38 3.34 1.13 97.58 2.03 0.39 94.24 1.81 0.73
NUCFL (DCA+COS) 94.50 0.45 0.20 93.47 0.71 0.35 93.35 3.43 1.25 97.68 1.17 0.37 95.35 0.61 0.22
NUCFL (MDCA+COS) 95.39 0.42 0.21 93.78 1.45 0.48 93.44 3.29 1.03 97.87 1.25 0.33 95.83 0.45 0.20
NUCFL (DCA+L-CKA) 94.91 0.32 0.18 93.32 0.53 0.31 92.98 2.84 0.98 97.80 1.11 0.36 95.80 0.40 0.14
NUCFL (MDCA+L-CKA) 95.36 0.39 0.21 93.09 0.41 0.27 92.67 3.09 1.04 97.83 0.97 0.27 94.86 0.59 0.21
NUCFL (DCA+RBF-CKA) 95.61 0.50 0.19 93.46 0.31 0.25 92.68 3.26 1.16 97.19 1.22 0.32 94.87 0.52 0.19
NUCFL (MDCA+RBF-CKA) 95.85 0.48 0.19 93.69 0.57 0.23 92.18 3.04 0.91 97.32 1.31 0.43 94.18 0.53 0.18
@@
Table 9: Accuracy (%), calibration measures ECE (%), and SCE (%) of various federated optimization methods with different calibration methods under non-IID ($\alpha=0.1$) scenario on the MNIST dataset. Values in boldface indicate the best calibration across all methods, achieved without compromising accuracy compared to the uncalibrated model.
@@
Calibration Method FedAvg McMahan et al. [2017] FedProx Sahu et al. [2018] Scaffold Karimireddy et al. [2019] FedDyn Acar et al. [2021] FedNova Wang et al. [2020b]
Uncal. 90.51 5.22 1.33 85.02 5.60 2.26 84.20 13.96 3.85 89.51 4.52 1.57 88.11 1.68 1.49
Focal Lin et al. [2017] 88.95 15.43 3.23 83.22 14.26 3.83 81.50 24.37 5.74 88.49 3.24 1.46 88.23 9.74 2.65
LS Müller et al. [2019] 88.48 50.62 10.57 77.52 34.49 9.32 75.89 41.50 10.87 84.95 6.20 4.66 80.84 26.31 7.87
BS Brier [1950] 84.36 12.65 2.90 74.96 9.44 4.14 69.21 18.55 7.03 83.05 5.85 3.87 78.29 2.79 3.34
MMCE Kumar et al. [2018] 93.99 13.18 2.74 83.32 16.74 3.58 77.48 29.33 6.48 88.49 3.48 1.16 89.48 9.90 2.34
FLSD Mukhoti et al. [2020] 87.42 29.42 5.78 81.52 28.11 6.46 77.73 37.03 8.76 86.86 5.87 1.81 87.00 21.39 4.76
MbLS Liu et al. [2021] 89.96 21.51 4.71 79.66 13.59 4.80 80.73 22.71 6.25 89.40 5.05 2.03 82.47 6.62 4.27
DCA Liang et al. [2020] 91.48 3.83 1.01 85.30 2.68 1.91 84.39 9.75 3.18 89.52 3.95 1.07 89.57 0.99 1.35
MDCA Hebbalaguppe et al. [2022b] 90.36 5.41 1.38 84.13 4.55 2.24 83.52 12.51 3.65 90.30 4.06 1.21 89.60 1.92 1.34
NUCFL (DCA+COS) 92.16 1.38 0.83 89.07 1.03 0.91 87.40 2.28 1.99 89.62 2.48 1.04 90.73 0.66 0.27
NUCFL (MDCA+COS) 92.20 1.30 0.74 89.29 1.89 1.10 86.94 3.37 1.80 90.90 2.72 0.79 90.76 1.25 0.30
NUCFL (DCA+L-CKA) 91.40 1.20 0.87 90.86 0.90 0.76 84.93 3.65 1.33 90.61 2.39 0.90 90.80 0.89 0.26
NUCFL (MDCA+L-CKA) 92.33 1.63 0.80 90.51 0.86 0.67 85.03 3.19 1.67 90.89 2.15 0.70 90.47 0.60 0.19
NUCFL (DCA+RBF-CKA) 91.06 1.38 0.78 90.31 0.59 0.58 84.93 3.14 1.76 90.13 2.49 0.84 89.35 0.71 0.23
NUCFL (MDCA+RBF-CKA) 92.15 1.77 0.81 90.00 0.71 0.65 85.71 3.79 2.11 90.76 2.18 0.73 90.73 0.55 0.19
@@
Table 10: Accuracy (%), calibration measures ECE (%), and SCE (%) of various federated optimization methods with different calibration methods under non-IID ($\alpha=0.05$) scenario on the MNIST dataset. Values in boldface indicate the best calibration across all methods, achieved without compromising accuracy compared to the uncalibrated model.

### v3

@@ B.2 Additional Results of Calibration on FL
Calibration Method FedAvg (McMahan et al., 2017) FedProx (Sahu et al., 2018) Scaffold (Karimireddy et al., 2019)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
Uncal. 92.68 $\pm$ 0.35 2.96 $\pm$ 0.22 1.04 $\pm$ 0.14 91.26 $\pm$ 0.63 4.21 $\pm$ 0.28 1.20 $\pm$ 0.11 92.00 $\pm$ 1.18 6.03 $\pm$ 0.37 1.92 $\pm$ 0.18
Focal (Lin et al., 2017) 90.25 $\pm$ 0.68 7.02 $\pm$ 0.27 2.58 $\pm$ 0.15 91.11 $\pm$ 0.95 6.93 $\pm$ 1.35 1.99 $\pm$ 0.55 88.24 $\pm$ 0.82 6.49 $\pm$ 1.10 1.92 $\pm$ 0.35
LS (Müller et al., 2019) 92.51 $\pm$ 0.28 4.28 $\pm$ 2.10 2.33 $\pm$ 0.48 90.26 $\pm$ 0.85 5.48 $\pm$ 1.45 1.62 $\pm$ 0.38 90.84 $\pm$ 1.25 9.26 $\pm$ 2.25 4.84 $\pm$ 0.50
BS (Brier, 1950) 90.09 $\pm$ 0.85 6.06 $\pm$ 0.32 2.49 $\pm$ 0.17 89.50 $\pm$ 0.48 5.58 $\pm$ 0.25 1.68 $\pm$ 0.75 90.06 $\pm$ 1.05 8.45 $\pm$ 0.55 3.03 $\pm$ 0.20
MMCE (Kumar et al., 2018) 90.22 $\pm$ 0.75 10.20 $\pm$ 0.35 5.16 $\pm$ 0.22 90.98 $\pm$ 0.70 8.28 $\pm$ 0.50 3.00 $\pm$ 0.17 89.87 $\pm$ 1.15 7.49 $\pm$ 0.28 2.61 $\pm$ 0.20
FLSD (Mukhoti et al., 2020) 90.57 $\pm$ 0.55 8.26 $\pm$ 0.24 3.01 $\pm$ 0.16 88.65 $\pm$ 0.75 6.00 $\pm$ 0.40 1.70 $\pm$ 0.20 90.37 $\pm$ 1.00 7.66 $\pm$ 0.45 2.65 $\pm$ 0.24
MbLS (Liu et al., 2021) 93.00 $\pm$ 1.20 5.22 $\pm$ 1.60 1.58 $\pm$ 0.38 91.47 $\pm$ 0.90 9.60 $\pm$ 0.70 4.86 $\pm$ 0.30 89.82 $\pm$ 0.75 9.29 $\pm$ 0.60 4.80 $\pm$ 0.22
DCA (Liang et al., 2020) 92.99 $\pm$ 0.35 2.06 $\pm$ 0.18 0.81 $\pm$ 0.22 91.95 $\pm$ 0.60 3.62 $\pm$ 0.45 1.04 $\pm$ 0.33 92.84 $\pm$ 0.80 4.49 $\pm$ 0.70 1.52 $\pm$ 0.40
NUCFL (DCA+COS) 93.00 $\pm$ 0.40 1.45 $\pm$ 0.20 0.70 $\pm$ 0.12 92.07 $\pm$ 0.48 2.81 $\pm$ 0.50 1.00 $\pm$ 0.18 92.88 $\pm$ 0.70 4.43 $\pm$ 0.65 1.52 $\pm$ 0.30
NUCFL (DCA+L-CKA) 92.41 $\pm$ 0.30 1.32 $\pm$ 0.15 0.66 $\pm$ 0.14 92.12 $\pm$ 0.55 2.83 $\pm$ 0.35 1.06 $\pm$ 0.20 92.44 $\pm$ 0.85 4.34 $\pm$ 0.55 1.48 $\pm$ 0.28
NUCFL (DCA+RBF-CKA) 93.11 $\pm$ 0.51 1.40 $\pm$ 0.22 0.71 $\pm$ 0.10 91.96 $\pm$ 0.65 2.83 $\pm$ 0.28 1.05 $\pm$ 0.15 92.18 $\pm$ 0.75 4.26 $\pm$ 0.70 1.40 $\pm$ 0.35
MDCA (Hebbalaguppe et al., 2022b) 92.94 $\pm$ 0.43 2.53 $\pm$ 0.35 0.89 $\pm$ 0.25 91.31 $\pm$ 0.70 3.78 $\pm$ 0.50 1.08 $\pm$ 0.28 91.88 $\pm$ 0.90 4.39 $\pm$ 0.65 1.53 $\pm$ 0.30
NUCFL (MDCA+COS) 92.89 $\pm$ 0.38 2.00 $\pm$ 0.18 0.82 $\pm$ 0.15 91.28 $\pm$ 0.60 3.05 $\pm$ 0.40 1.03 $\pm$ 0.22 92.04 $\pm$ 0.80 4.22 $\pm$ 0.60 1.50 $\pm$ 0.28
NUCFL (MDCA+L-CKA) 93.86 $\pm$ 0.33 1.95 $\pm$ 0.20 0.81 $\pm$ 0.12 92.09 $\pm$ 0.75 3.40 $\pm$ 0.30 1.05 $\pm$ 0.18 92.17 $\pm$ 0.85 4.13 $\pm$ 0.50 1.37 $\pm$ 0.25
NUCFL (MDCA+RBF-CKA) 93.35 $\pm$ 0.40 1.88 $\pm$ 0.25 0.79 $\pm$ 0.14 92.19 $\pm$ 0.55 3.57 $\pm$ 0.35 1.08 $\pm$ 0.17 91.68 $\pm$ 0.95 4.13 $\pm$ 0.60 1.40 $\pm$ 0.28
Calibration Method FedDyn (Acar et al., 2021) FedNova (Wang et al., 2020b)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
Uncal. 94.24 $\pm$ 0.90 4.28 $\pm$ 0.72 1.08 $\pm$ 0.09 94.30 $\pm$ 1.12 2.95 $\pm$ 0.35 0.89 $\pm$ 0.13
Focal (Lin et al., 2017) 95.03 $\pm$ 0.45 7.38 $\pm$ 0.65 2.61 $\pm$ 0.11 93.00 $\pm$ 1.15 9.32 $\pm$ 0.55 4.82 $\pm$ 0.13
LS (Müller et al., 2019) 95.16 $\pm$ 1.05 10.86 $\pm$ 1.65 5.20 $\pm$ 0.30 93.15 $\pm$ 1.18 9.00 $\pm$ 1.70 4.80 $\pm$ 0.95
BS (Brier, 1950) 95.26 $\pm$ 0.70 9.20 $\pm$ 0.47 4.83 $\pm$ 0.13 89.59 $\pm$ 0.95 4.23 $\pm$ 0.30 1.19 $\pm$ 0.16
MMCE (Kumar et al., 2018) 93.36 $\pm$ 0.60 5.95 $\pm$ 0.38 1.74 $\pm$ 0.14 93.88 $\pm$ 1.05 8.89 $\pm$ 0.55 4.77 $\pm$ 0.22
FLSD (Mukhoti et al., 2020) 93.37 $\pm$ 0.50 5.67 $\pm$ 0.43 1.70 $\pm$ 0.12 92.13 $\pm$ 0.85 10.97 $\pm$ 0.65 5.21 $\pm$ 0.27
MbLS (Liu et al., 2021) 91.52 $\pm$ 0.85 4.16 $\pm$ 0.50 1.06 $\pm$ 0.14 94.84 $\pm$ 1.30 4.21 $\pm$ 0.75 1.12 $\pm$ 0.25
DCA (Liang et al., 2020) 95.25 $\pm$ 0.50 3.18 $\pm$ 0.28 1.02 $\pm$ 0.15 93.84 $\pm$ 0.95 1.87 $\pm$ 0.38 0.79 $\pm$ 0.18
NUCFL (DCA+COS) 95.08 $\pm$ 0.65 2.57 $\pm$ 0.33 0.87 $\pm$ 0.22 94.85 $\pm$ 0.80 1.61 $\pm$ 0.28 0.75 $\pm$ 0.14
NUCFL (DCA+L-CKA) 94.30 $\pm$ 0.45 2.41 $\pm$ 0.25 0.86 $\pm$ 0.17 94.00 $\pm$ 0.90 1.40 $\pm$ 0.22 0.70 $\pm$ 0.12
NUCFL (DCA+RBF-CKA) 95.69 $\pm$ 0.55 2.50 $\pm$ 0.30 0.89 $\pm$ 0.14 93.37 $\pm$ 0.85 1.52 $\pm$ 0.25 0.73 $\pm$ 0.13
MDCA (Hebbalaguppe et al., 2022b) 95.08 $\pm$ 0.65 3.13 $\pm$ 0.38 1.02 $\pm$ 0.20 93.74 $\pm$ 1.00 2.77 $\pm$ 0.42 0.86 $\pm$ 0.24
NUCFL (MDCA+COS) 94.37 $\pm$ 0.55 2.85 $\pm$ 0.35 1.00 $\pm$ 0.18 95.23 $\pm$ 0.90 2.45 $\pm$ 0.30 0.82 $\pm$ 0.13
NUCFL (MDCA+L-CKA) 94.33 $\pm$ 0.45 2.84 $\pm$ 0.28 1.01 $\pm$ 0.15 94.36 $\pm$ 0.95 2.39 $\pm$ 0.25 0.80 $\pm$ 0.14
NUCFL (MDCA+RBF-CKA) 94.82 $\pm$ 0.60 2.51 $\pm$ 0.40 0.83 $\pm$ 0.22 93.68 $\pm$ 1.05 2.43 $\pm$ 0.28 0.80 $\pm$ 0.12
@@
Table 9: Accuracy (%), calibration measures ECE (%), and SCE (%) of various federated optimization methods with different calibration methods under non-IID ($\alpha=0.1$) scenario on the MNIST dataset. Underlined values indicate the best calibration across all methods.
@@
Calibration Method FedAvg (McMahan et al., 2017) FedProx (Sahu et al., 2018) Scaffold (Karimireddy et al., 2019)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
Uncal. 87.51 $\pm$ 0.45 6.22 $\pm$ 0.22 1.83 $\pm$ 0.15 87.02 $\pm$ 0.60 6.60 $\pm$ 0.30 2.22 $\pm$ 0.25 88.20 $\pm$ 0.80 10.96 $\pm$ 0.50 3.53 $\pm$ 0.35
Focal (Lin et al., 2017) 87.95 $\pm$ 1.55 10.43 $\pm$ 0.35 3.50 $\pm$ 0.20 84.22 $\pm$ 0.70 11.26 $\pm$ 0.95 3.73 $\pm$ 0.25 83.50 $\pm$ 0.85 8.37 $\pm$ 1.65 3.04 $\pm$ 0.88
LS (Müller et al., 2019) 86.48 $\pm$ 1.91 11.62 $\pm$ 1.15 3.91 $\pm$ 0.92 83.52 $\pm$ 0.69 7.49 $\pm$ 1.77 2.68 $\pm$ 0.88 84.89 $\pm$ 1.11 14.57 $\pm$ 1.85 4.64 $\pm$ 0.79
BS (Brier, 1950) 83.36 $\pm$ 0.85 5.65 $\pm$ 0.56 1.72 $\pm$ 0.35 87.96 $\pm$ 1.15 10.44 $\pm$ 1.61 3.64 $\pm$ 0.99 88.21 $\pm$ 1.55 19.55 $\pm$ 2.92 7.53 $\pm$ 1.05
MMCE (Kumar et al., 2018) 88.99 $\pm$ 1.65 12.18 $\pm$ 1.43 4.04 $\pm$ 0.77 86.32 $\pm$ 1.83 11.74 $\pm$ 0.96 3.75 $\pm$ 0.72 86.48 $\pm$ 1.29 14.33 $\pm$ 2.38 4.61 $\pm$ 0.67
FLSD (Mukhoti et al., 2020) 85.42 $\pm$ 0.77 9.28 $\pm$ 1.06 3.48 $\pm$ 0.45 85.52 $\pm$ 1.05 10.11 $\pm$ 1.85 3.60 $\pm$ 0.55 86.73 $\pm$ 1.35 11.03 $\pm$ 1.11 3.69 $\pm$ 0.78
MbLS (Liu et al., 2021) 68.96 $\pm$ 1.81 13.27 $\pm$ 1.55 4.51 $\pm$ 0.40 87.66 $\pm$ 1.19 11.59 $\pm$ 1.57 3.75 $\pm$ 0.89 89.73 $\pm$ 2.25 18.66 $\pm$ 2.79 7.05 $\pm$ 1.25
DCA (Liang et al., 2020) 88.48 $\pm$ 1.15 5.83 $\pm$ 0.80 1.75 $\pm$ 0.22 87.30 $\pm$ 0.99 5.68 $\pm$ 0.35 1.72 $\pm$ 0.28 88.69 $\pm$ 1.01 9.22 $\pm$ 0.72 3.44 $\pm$ 0.37
NUCFL (DCA+COS) 88.00 $\pm$ 0.71 5.38 $\pm$ 0.25 1.69 $\pm$ 0.28 88.00 $\pm$ 1.11 5.03 $\pm$ 0.22 1.41 $\pm$ 0.77 88.82 $\pm$ 1.33 8.69 $\pm$ 0.77 3.29 $\pm$ 0.37
NUCFL (DCA+L-CKA) 87.70 $\pm$ 1.14 5.20 $\pm$ 0.72 1.59 $\pm$ 0.22 87.33 $\pm$ 1.05 5.05 $\pm$ 0.17 1.41 $\pm$ 0.24 88.71 $\pm$ 1.13 8.55 $\pm$ 0.35 3.23 $\pm$ 0.20
NUCFL (DCA+RBF-CKA) 88.79 $\pm$ 1.03 5.20 $\pm$ 1.22 1.61 $\pm$ 0.87 88.14 $\pm$ 1.19 5.11 $\pm$ 0.20 1.47 $\pm$ 0.79 89.01 $\pm$ 1.03 8.52 $\pm$ 0.61 3.23 $\pm$ 0.61
MDCA (Hebbalaguppe et al., 2022b) 88.36 $\pm$ 1.60 5.41 $\pm$ 0.79 1.72 $\pm$ 0.25 88.13 $\pm$ 1.95 5.84 $\pm$ 0.81 1.77 $\pm$ 0.67 88.52 $\pm$ 1.15 9.51 $\pm$ 0.75 3.47 $\pm$ 0.30
NUCFL (MDCA+COS) 88.20 $\pm$ 1.21 5.30 $\pm$ 0.71 1.67 $\pm$ 0.29 88.29 $\pm$ 1.21 5.21 $\pm$ 0.67 1.60 $\pm$ 0.29 88.64 $\pm$ 1.09 8.67 $\pm$ 0.82 3.28 $\pm$ 0.29
NUCFL (MDCA+L-CKA) 87.33 $\pm$ 1.03 5.33 $\pm$ 0.64 1.68 $\pm$ 0.39 87.51 $\pm$ 1.40 5.11 $\pm$ 0.32 1.49 $\pm$ 0.09 88.60 $\pm$ 1.10 8.52 $\pm$ 0.35 3.22 $\pm$ 0.22
NUCFL (MDCA+RBF-CKA) 88.15 $\pm$ 0.96 5.27 $\pm$ 0.32 1.66 $\pm$ 0.15 88.00 $\pm$ 1.03 5.38 $\pm$ 0.20 1.70 $\pm$ 0.12 88.43 $\pm$ 1.00 8.69 $\pm$ 0.61 3.27 $\pm$ 0.18
Calibration Method FedDyn (Acar et al., 2021) FedNova (Wang et al., 2020b)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
Uncal. 88.51 $\pm$ 0.70 7.52 $\pm$ 0.40 2.67 $\pm$ 0.28 89.11 $\pm$ 0.90 9.68 $\pm$ 0.18 3.49 $\pm$ 0.22
Focal (Lin et al., 2017) 87.49 $\pm$ 1.65 9.24 $\pm$ 0.77 3.46 $\pm$ 0.93 87.23 $\pm$ 0.81 10.74 $\pm$ 0.50 3.50 $\pm$ 0.27
LS (Müller et al., 2019) 86.95 $\pm$ 1.85 7.20 $\pm$ 0.38 2.65 $\pm$ 0.32 85.84 $\pm$ 1.20 8.31 $\pm$ 0.77 3.27 $\pm$ 0.92
BS (Brier, 1950) 84.05 $\pm$ 0.95 6.85 $\pm$ 0.55 2.35 $\pm$ 0.33 87.29 $\pm$ 1.70 9.79 $\pm$ 1.28 3.49 $\pm$ 0.44
MMCE (Kumar et al., 2018) 87.49 $\pm$ 1.74 9.48 $\pm$ 1.22 3.50 $\pm$ 0.92 88.48 $\pm$ 1.10 11.90 $\pm$ 0.65 3.84 $\pm$ 0.69
FLSD (Mukhoti et al., 2020) 85.86 $\pm$ 1.10 6.97 $\pm$ 0.44 2.37 $\pm$ 0.31 86.00 $\pm$ 1.39 10.39 $\pm$ 2.80 3.46 $\pm$ 1.50
MbLS (Liu et al., 2021) 88.40 $\pm$ 0.95 9.05 $\pm$ 1.40 3.43 $\pm$ 0.92 87.47 $\pm$ 1.43 10.62 $\pm$ 1.55 3.51 $\pm$ 0.82
DCA (Liang et al., 2020) 88.52 $\pm$ 0.83 6.95 $\pm$ 0.39 2.37 $\pm$ 0.25 88.97 $\pm$ 1.00 8.09 $\pm$ 0.39 3.09 $\pm$ 0.13
NUCFL (DCA+COS) 89.00 $\pm$ 0.99 6.48 $\pm$ 0.28 2.30 $\pm$ 0.18 89.29 $\pm$ 1.11 7.66 $\pm$ 0.20 2.66 $\pm$ 0.73
NUCFL (DCA+L-CKA) 88.71 $\pm$ 1.23 6.39 $\pm$ 1.03 2.27 $\pm$ 0.82 89.50 $\pm$ 0.61 7.79 $\pm$ 1.02 2.72 $\pm$ 0.68
NUCFL (DCA+RBF-CKA) 88.90 $\pm$ 0.99 6.23 $\pm$ 0.73 2.25 $\pm$ 0.48 88.92 $\pm$ 0.61 6.97 $\pm$ 0.72 2.45 $\pm$ 0.17
MDCA (Hebbalaguppe et al., 2022b) 89.03 $\pm$ 1.25 6.26 $\pm$ 0.77 2.21 $\pm$ 0.49 88.60 $\pm$ 1.23 8.12 $\pm$ 1.05 3.09 $\pm$ 0.49
NUCFL (MDCA+COS) 88.90 $\pm$ 1.29 5.91 $\pm$ 0.68 1.79 $\pm$ 0.39 89.76 $\pm$ 1.29 7.61 $\pm$ 0.68 2.67 $\pm$ 0.23
NUCFL (MDCA+L-CKA) 89.09 $\pm$ 1.24 5.88 $\pm$ 0.49 1.79 $\pm$ 0.18 90.07 $\pm$ 1.15 7.55 $\pm$ 0.66 2.64 $\pm$ 0.37
NUCFL (MDCA+RBF-CKA) 88.76 $\pm$ 0.67 5.88 $\pm$ 1.28 1.80 $\pm$ 0.44 89.73 $\pm$ 1.33 7.41 $\pm$ 0.76 2.60 $\pm$ 0.09
@@
Table 10: Accuracy (%), calibration measures ECE (%), and SCE (%) of various federated optimization methods with different calibration methods under non-IID ($\alpha=0.05$) scenario on the MNIST dataset. Underlined values indicate the best calibration across all methods.